OpenStudy (anonymous):

Find the equation that passes through (-3,-3) and is parallel to x-2y= -7

OpenStudy (gorv):

first convert given equation in standard form y =mx+c x+7=2y 2y=x+7 y=x/2+7/2

OpenStudy (anonymous):

Beat me to it gorv xD

OpenStudy (gorv):

slope of given line =m=1/2 parallel line have same slope so slope of parallel line =1/2

OpenStudy (gorv):

line pass through (-3,-3)

OpenStudy (gorv):

put all in y=mx+c -3=1/2*(-3)+c -3=-3/2+c c=-3+3/2=-3/2

OpenStudy (gorv):

so equation is y=x/2-3/2